# Unveiling the Siren’s Song: Towards Reliable Fact-Conflicting Hallucination Detection

### v1

@@ 6. Conclusion and Future Work
We introduce FactCHD, a meticulously designed benchmark tailored for the rigorous evaluation of fact-conflicting hallucinations within LLMs, which is notably enriched with a kaleidoscope of patterns and substantiated chain of evidence to fortify the robust elucidation of factuality assessments. Additionally, we unveil a semi-supervised data construction strategy, leveraging knowledge as facts to facilitate the creation of a hallucination dataset. Moreover, we delineate Truth-Triangulator, an innovative framework that elegantly employs the principle of triangulation to astutely discern the veracity of information, deploying cross-referencing generators and search tools to judiciously arbitrate responses, especially those teetering on the brink of questionable factuality.
Moving forward, our future endeavors will focus on three primary objectives: 1) explore efficient and powerful knowledge enhancement methods to enhance hallucination; 2) investigate factual errors stemming from incorrect or outdated prior knowledge, as well as the potential occurrence of hallucination during inference; 3) broadening our evaluative scope to encompass various modalities and cultural contexts in the assessment of hallucinations.
@@ Appendix A In-depth Exploration into Benchmark Construction
Navigating through the space constraints, this section delicately provides a condensed, yet comprehensive overview of the data collection process, illustrated meticulously in Figure 6. We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 6, 7, 8, and 9. Further, the meticulous design and specific prompt crucial for the chain of evidence generation are depicted in Table 10, offering a peek into the intricate mechanisms behind our methodology.
Appendix B A Glimpse into Experimental Intricacies
@@
Figure 6. Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
@@
Table 6. Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
Table 7. Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
Table 8. Instruction for generating the “Query-Response” context based on KG with the comparing fact. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
Table 9. Instruction for generating the “Query-Response” context based on KG with the set operation. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
Table 10. Instruction to generate evidence-based chains of facts for the“Query-Response” contexts. The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
I want you to act as a data generator for a fallacy find task. You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#. I’ll tell you the right label: FACTUAL or NON-FACTUAL. Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#. Here are some examples:
#Golden Label#: FACTUAL
#Query#: Can you tell me which country is Chris Terrios from?
#Response#: Chris Terrio is an American.
#Evidence#: [Chris_Terrio: Chris Terrio (born December 31, 1976) is an American screenwriter and film director.]
#Output#: FACTUAL. The answer that Chris Terrio is an American is correct. Chris Terrio is an American screenwriter and film director. So there are no fallacies, faulty reasoning, or incorrect conclusions present in this example.
#Golden Label#: NON-FACTUAL
#Query#: How many children did the head of government of the country where the distributor of "The Killer Inside Me" is located have?
#Golden Lable#: <Here is the right label for the factuality of the following sample >
#Query#: <Here is the query for the LLM>
#Response#: <Here is the response from the LLM according to the above query>
#Evidence#: <Here is the evidence with fact to judge the factuality of above context>
#Output#:
@@
Table 11. Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
@@
Table 12. Prompts for few-shot evaluation. Note that the demonstration only adopted in ICL(4-shot) setting. Large language models are known to be highly sensitive to prompts, especially in zero-shot and in-context learning settings. Therefore, we cannot guarantee that the prompts we used in the evaluation of FACTLLM are optimal. The role description is indicated by the yellow text, the evaluation demonstration is represented by the purple text.
@@
Query: Could you tell me the name of the series that Rick Yune was on that got cancelled on December 12?
Response: Rick Yune was on a series that cancelled on December 12.
@@
FACTUAL. The answer that Rick Yune was on a series that got cancelled on December 12 is correct. Rick Yune was a cast member of the television series The Blacklist, which was cancelled on December 12, 2014. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
@@
Table 13. Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.

### v2

@@ 6 Conclusion and Future Work
We introduce FactCHD, a meticulously designed benchmark tailored for the evaluation of fact-conflicting hallucinations from LLMs, which is notably enriched with a kaleidoscope of patterns and substantiated evidence chains to fortify the robust elucidation of factuality assessments. Moreover, we delineate Truth-Triangulator that employs the principle of triangulation to discern the veracity of information, deploying cross-referencing generators to arbitrate responses. Moving forward, we will broaden our evaluative scope to encompass various modalities and granularity in hallucination detection.
@@ Appendix A In-depth Exploration into Benchmark Construction
Navigating through the space constraints, this section delicately provides a condensed, yet comprehensive overview of the data collection process, illustrated meticulously in Figure 7. We shed light on the intriguing prompt of ”Query-Response” generation, which seamlessly intertwines with four distinct actuality patterns, brought to life in Tables 7, 8, 9, and 10.
Appendix B Experimental Details
@@
Figure 7: Knowledge as fact for constructing the benchmark on detecting fact-conflicting hallucination.
@@
Table 7: Instruction for generating the “Query-Response” context based on FEVER-like dataset with the vanilla factuality pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
Table 8: Instruction for generating the “Query-Response” context based on KG with the multi-hop reasoning pattern. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
Table 9: Instruction for generating the “Query-Response” context based on KG with the comparing fact. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
Table 10: Instruction for generating the “Query-Response” context based on KG with the set operation. The role description is indicated by the yellow text, the factuality demonstration is represented by the purple text, while the factuality pattern is descripited by the blue text.
@@
Table 11: Prompts for zero-shot evaluation. The role description is indicated by the yellow text.
@@
Table 12: Prompts for few-shot evaluation. Note that the demonstration only adopted in ICL(4-shot) setting. Large language models are known to be highly sensitive to prompts, especially in zero-shot and in-context learning settings. Therefore, we cannot guarantee that the prompts we used in the evaluation of FACTLLM are optimal. The role description is indicated by the yellow text, the evaluation demonstration is represented by the purple text.
@@
Query: Could you tell me the name of the series that Rick Yune was on that got canceled on December 12?
Response: Rick Yune was on a series that canceled on December 12.
@@
FACTUAL. The answer that Rick Yune was on a series that got canceled on December 12 is correct. Rick Yune was a cast member of the television series The Blacklist, which was canceled on December 12, 2014. Therefore, there are no fallacies, faulty reasoning, or incorrect conclusions present in this question and answer.
@@
Table 13: Case example from the results of Alpaca-7B-LoRa (Tuning) and the illustration of the impact of retrieval enhancement.